### v2

# PriorDiffusion: Leverage Language Prior in Diffusion Models for Monocular Depth Estimation

Abstract
Traditional monocular depth estimation suffers from inherent ambiguity and visual nuisance. We argue that language prior can enhance monocular depth estimation by leveraging the inductive bias learned during the text-to-image pre-training of diffusion models. The ability of these models to generate images that align with text indicates that they have learned the spatial relationships, size, and shape of specified objects, which can be applied to improve depth estimation. Thus, we propose PriorDiffusion, using a pre-trained text-to-image diffusion model that takes both images and corresponding text descriptions to infer affine-invariant depth through a denoising process. We also show that language prior enhances the model’s perception of specific regions of images that users care about and describe. Simultaneously, language prior acts as a constraint to accelerate the convergence of both training and the inference diffusion trajectory. By training on HyperSim and Virtual KITTI, we achieve faster training convergence, fewer inference diffusion steps, and state-of-the-art zero-shot performance across NYUv2, KITTI, ETH3D, and ScanNet. Code will be released upon acceptance.
Figure 1: Language prior in diffusion models enhances depth estimation by leveraging the geometric prior aligned with the language description. It is learned during text-to-image generation pre-training, as to generate images under different viewpoints and layouts that accurately reflect the text, the model must comprehend the size and shape of specified objects, their spatial relationship, and the scale of the scene.
1 Introduction
Monocular depth estimation requires the model to predict pixel-wise depth from a single image. Recent advancements in diffusion-based models have shown promise for generating high-quality depth maps [12, 37, 69, 70, 110, 38, 19, 29, 17, 23, 31, 87] due to their ability to capture complex structures and fine details by progressively denoising a latent representation to produce accurate depth predictions. However, it remains a challenging task due to inherent ambiguities and various visual nuisances. Unlike stereo or multi-view depth estimation methods, monocular approaches lack depth cues provided by multiple perspectives, leading to difficulty in accurately discerning the relative sizes, distances, and spatial relationships of objects within a scene. That is, different real-world 3D scenes can project to the same 2D image, causing inherent ambiguities. For example, texture ambiguity. Surfaces that are uniformly textured or have repetitive patterns, like tiled floors, can create confusion for depth estimators. The models might misinterpret the repetition as uniform depth, even though the actual depth might vary (e.g., a tiled floor receding into the distance). Also, visibility ambiguity creates limitations for discerning small, partial or occluded objects purely from the visual signal. On the other hand, visual nuisances such as illumination, motion blur, atmospheric conditions, and perspective distortion distort or obscure critical visual cues, leading vision algorithms to misinterpret scene structure, depth, and object boundaries, thus reducing their accuracy and reliability.
Shown in Figure 1, an emerging solution lies in leveraging language as a complementary modality to provide a prior to resolve ambiguity and visual nuisance. Language descriptions provided by humans offer a wealth of priors about spatial relationships, scene types, object sizes, and depth hierarchies.
In cases where an object is small, partially occluded, visually similar to the background, or its type is ambiguous from a single image, language can indicate the object’s identity and semantics, enabling the model to infer its typical 3D size and shape and make more precise depth predictions for the observed portion. Language inputs are also valuable in resolving texture ambiguities. Descriptive cues can specify the type and function of surfaces in a scene, aiding the model in inferring geometric properties. For instance, if a uniformly textured surface is described as a “white ceiling”, the model can interpret it as a continuous, flat plane at a consistent depth above the observer, extending into the distance. Similarly, for surfaces with repetitive patterns, such as a red and black interwoven carpet, language inputs like “A red and black interwoven carpet extending across the room” can help the model recognize that the repeating pattern belongs to a flat surface, preventing misinterpretation of edges between tiles as depth discontinuities. As shown in Figure 2, better perception can be achieved for insignificant regions indicated in the language description.
To this end, we propose PriorDiffusion, incorporating human-provided language prior of the scene as an inductive bias to guide the depth map prediction. During text-to-image pre-training, diffusion models learn to generate diverse images under various viewpoints and scene layouts that align with the provided language descriptions. The ability of these models to generate images that align with text suggests they have learned the spatial relationships, size, and shape of specified objects in the described 3D scene, which can be leveraged to improve depth estimation. In our PriorDiffusion, during the denoising process, both the image and the text input are used by the model as conditions to predict the noise to be removed. Ultimately, Gaussian noise is progressively refined into a depth map that aligns with both the input image and the language description. In real-world applications when embodied agents are integrated with PriorDiffusion, humans can describe the environment to help agents better perceive it. When generalizing to a new environment, language, having less covariate shift compared to images, allows humans to describe the new environment, aiding agents in better adapting and generalizing to it.
To support our hypothesis, we conduct training on two synthetic datasets, HyperSim [65] and Virtual KITTI [4], and conduct zero-shot evaluation on four real-world datasets, NYUv2 [72], KITTI [25], ETH3D [71], and ScanNet [10]. Given the difficulty of obtaining sufficient human-provided text descriptions for every training and inference image, we utilize a vision-language model, i.e. LLaVA [11], to generate descriptions for each image, simulating human annotation. we achieve faster training convergence, fewer inference diffusion steps, and state-of-the-art zero-shot performance across NYUv2, KITTI, ETH3D, and ScanNet.
Our contributions. By leveraging language prior for diffusion-based monocular depth estimation, we demonstrate that:
  • Language prior provides an inductive bias about the existence, geometric properties, and spatial relationships of objects specified in the language description, allowing depth estimators to better perceive the depth of those objects.
  • Language prior directs the model’s attention to specific regions, guiding it to perceive the 3D scene according to the user’s intent. This is particularly beneficial for regions that pose challenges to vision systems, such as those with small size, poor illumination, occlusion, or high visual similarity to the background.
  • Language prior serves as a constraint to speed up training convergence, and serves as a good initialization of diffusion trajectory to speed up inference. This improves the efficiency of the diffusion model, optimizing both its performance and computational resource usage.
Figure 2: 3D reconstruction. For insignificant regions indicated in language description, PriorDiffusion achieves better depth estimates and reconstructions thanks to the use of language.
Figure 3: Pipeline of PriorDiffusion. We train the diffusion model to predict the noise added into the noisy depth latent $\mathbf{z}_{t}$ at the time step $t$, based on $\mathbf{z}_{t}$, the input image $x$, and the corresponding language description $c$. During inference, the diffusion model predicts noise for $\mathbf{z}_{t}$ at each time step and gradually denoise it from $\mathbf{z}_{T}$ (pure Gaussian noise) into $\mathbf{z}_{0}$ (pure depth latent). Then $\mathbf{z}_{0}$ is decoded into the depth prediction using a frozen variational decoder.
2 Related Work
Monocular depth estimation. Monocular depth estimation is a dense regression problem, requiring pixel-wise predictions of depth from a single image, capturing complex spatial relationships within the scene. Some depth models learn to infer pixel-wise depth in metric scale (i.e. meters) [2, 18, 39, 48, 81, 97, 106, 98] by minimizing loss between depth predictions and ground-truth depth maps. Each model typically applies to only one data domain in which it is trained, with similar camera parameters and object scales.
When ground-truth depth is not available, self-supervised approaches [3, 17, 36, 40, 41, 53, 57, 75, 77, 78, 80, 85, 96, 102, 107, 108, 112, 115] rely on geometric constraints, where scale is attributed through radar [73], inertials [30, 79], lidar [16, 51, 56, 84, 83, 82, 86, 93], or binocular images [24, 28, 27]. To enable generalization across diverse domains, empowering depth estimation in the wild, other works have explored affine-invariant depth estimation [100, 90, 63, 62, 38, 91, 95, 13, 94, 47], which predicts depth up to an unknown global shift and scale. This approach can accommodate varied scenes with different scales, while still preserving the geometric relationships between different objects or regions within the scene. HND [100] hierarchically normalizes the depth representations with spatial information and depth distributions. Depth Anything  [90] learns from large-scale automatically annotated data. DPT [63] leverages vision transformers using a scale- and shift-invariant trimmed loss. MiDas [62] mixes multiple datasets with training objectives invariant to depth range and scale. Depth Anything [90] and Depth Anything v2 [91] train a depth foundation model using large-scale unlabeled data. Existing methods try to predict depth solely from images, which suffers from ambiguity like scale, object orientation, occlusion, and visual nuisance like viewpoints, illumination, appearance, texture, etc. On the other hand, our PriorDiffusion leverages language prior in Diffusion Models to provide a thorough understanding of the scene associated with the input text, to resolve the ambiguity and the nuisance variables in the visual algorithm.
Diffusion model for monocular depth estimation. Denoising Diffusion Probabilistic Model (DDPM) [33] learns to reverse a diffusion process that progressively degrades images with Gaussian noise so that they can draw samples from the data distribution by applying the reverse process to random noise. They have been applied to various tasks like text-to-image generation [66, 61, 119, 118, 76], super resolution [67, 44, 22], image inpainting [52, 9, 92], 3D object generation [58, 43, 50], etc. Several approaches have explored the use of DDPM for metric depth estimation. DiffusionDepth [12] learns to denoise random depth distribution into a depth map with monocular visual conditions. Similarly, DDP [37] encodes the input image and use diffusion to decodes it into a depth map. DepthGen [69] extends a multi-task diffusion model for metric depth prediction. Its successor, DDVM [70], focuses on pretraining with synthetic and real datasets to improve depth estimation performance. Build upon the diffusion model, DepthFM [29] uses flow matching to conduct a faster monocular depth estimation. In light of strong representation learned by pre-trained, text-to-image diffusion models, recent methods like VPD [110], E2E FT [23], GeoWizard [19], Lotus [31], GenPercept [87], and Marigold [38] utilize a pretrained Stable Diffusion model [66] as a backbone to predict depth through a denoising process. Specifically, Marigold [38] concatenates input images with diffusion latent as conditions for denoising U-Net to predict noise that needs to be removed, to gradually denoise a Gaussian noise to a depth map. While previous works only focused on using images as conditions of diffusion models to predict depth, our PriorDiffusion utilizes language prior as inductive bias for diffusion models to resolve ambiguity and visual nuisances, and conduct 3D perception in a controllable way.
Figure 4: Visualization on NYUv2. Compared to Margold, our PriorDiffusion demonstrates more accurate depth prediction for a given input image, particularly for instances specified in the language description (marked in red). This is achieved by leveraging the language prior within the text-to-image diffusion model, enhancing the model’s ability to comprehend the geometric characteristics of these objects. Additionally, the language description directs the model’s attention to relevant regions, especially the ones that are easily neglected by visual signals, like “a soap dispenser” in the first row. It enables the depth estimator to perceive the 3D scene in a more controlled manner that aligns with the user’s interest.
Language for monocular depth estimation. Vision-Language models  [5, 46, 45, 55, 60] acquire a comprehensive understanding of languages and images through pre-training under diverse datasets, thereby establishing a robust foundation for downstream tasks. [88, 89, 104, 7, 105, 106, 117, 98, 99, 11]. Typically, CLIP [60] employs contrastive learning on text-image pairs, enabling a range of applications like few-shot image classification  [21, 103, 59, 114], image segmentation  [64, 111], object detection  [64, 116], and 3D perception [35, 105, 106, 117, 113, 101, 8, 32]. In light of their emerging ability, some works [1, 35, 106, 109] have tried to apply vision-language models for monocular depth estimation. Particularly, DepthCLIP [106] leverages the semantic depth response of CLIP with a depth projection scheme to conduct zero-shot monocular depth estimation. WorDepth [98] learns the variational prior distribution of 3D scenes from text. RSA [99] predicts scale using language description to align relative depth to metric scale. While language has been applied in various aspects of 3D perception, our work is the first to use language prior for diffusion models to enhance 3D perception.
3 Method
Problem setup. Given an RGB image $x:\Omega\subset\mathbb{R}^{2}\rightarrow\mathbb{R}^{3}$, the goal of monocular depth estimation is to predict a dense depth map $y:\Omega\subset\mathbb{R}^{2}\rightarrow\mathbb{R}_{+}$ using a parameterized function $h$, represented as a neural network, such that $y:=h(\cdot)$. We utilize a supervised dataset $\mathcal{D}=\{x^{(m)},\textbf{c}^{(m)},y^{*(m)}\}_{m=1}^{M}$ containing $M$ samples, where $y^{*}:\Omega\subset\mathbb{R}^{2}\rightarrow\mathbb{R}_{+}$ is the ground-truth depth map, and c represents the text caption that describes the image.
Diffusion formulation. The pipeline of PriorDiffusion is shown in Figure 3. To use a diffusion model for monocular depth estimation, following [12, 37, 69, 70, 110, 38], we adopt the formulation of Denoising Diffusion Probabilistic Model (DDPM) [33]. We define a latent variable $\mathbf{z}_{t}$ $\in\mathbb{R}^{H^{\prime}\times W^{\prime}\times C^{\prime}}$that represents a noisy version of the depth map in the latent space at diffusion timestep $t$. The ground truch depth map $y^{*}$ is encoded into a latent representation $\mathbf{z_{y}}$ as supervision using a frozen VAE encoder $E$: $\mathbf{z_{y}}=E(y^{*})$. The diffusion latent after the whole denoising process $\mathbf{z_{0}}$ is decoded back to the image space using the frozen VAE decoder $D$: $\hat{y}=D(\mathbf{z}_{0})$.
Figure 5: Visualization on KITTI. PriorDiffusion predicts better depth for described objects, even when parts of the object are almost visible in the image (such as the parking car in the first column). The model is guided for specific described regions, such as a sign at a distance, potentially enhancing the safety of self-driving systems that rely solely on vision sensors.
Forward diffusion process. The forward process gradually adds noise to the latent depth feature in $T$ steps. The process can be defined as:
$q(\mathbf{z}_{t}|\mathbf{z}_{t-1})=\mathcal{N}(\mathbf{z}_{t};\sqrt{1-\beta_{t}}\mathbf{z}_{t-1},\beta_{t}\mathbf{I}),$
where $\beta_{t}$ is a scalar that defines the variance of the Gaussian noise added at each step $t$ in the forward diffusion process. This parameter controls how much noise is added into the latent variable $\mathbf{z}_{t}$ as it transitions from $\mathbf{z}_{t-1}$ over time.
The forward process from $y$ to $\mathbf{z}_{T}$ can be written as:
$q(\mathbf{z}_{1:T}|y^{*})=\prod_{t=1}^{T}q(\mathbf{z}_{t}|\mathbf{z}_{t-1}),$
where $D(\mathbf{z}_{0})=\hat{y}$.
Reverse diffusion process. The goal of the reverse process is to denoise $\mathbf{z}_{T}$ to recover $\hat{y}$. The reverse process is defined as:
$p_{\theta}(\mathbf{z}_{0:T}|x,\mathbf{c})=p(\mathbf{z}_{T})\prod_{t=1}^{T}p_{\theta}(\mathbf{z}_{t-1}|\mathbf{z}_{t},x,\mathbf{c}),$
where $p(\mathbf{z}_{T})$ is assumed to be a standard Gaussian distribution $\mathcal{N}(\mathbf{z}_{T};\mathbf{0},\mathbf{I})$, and the denoising step $p_{\theta}(\mathbf{z}_{t-1}|\mathbf{z}_{t},x,\mathbf{c})$ is parameterized as:
$p_{\theta}(\mathbf{z}_{t-1}|\mathbf{z}_{t})=\mathcal{N}(\mathbf{z}_{t-1};\mu_{\theta}(\mathbf{z}_{t},t,x,\mathbf{c}),\Sigma_{\theta}(\mathbf{z}_{t},t,x,\mathbf{c})),$
where $\mu_{\theta}$ and $\Sigma_{\theta}$ is the diffusion model, that predicts the mean and variance conditioned on $\mathbf{z}_{t}$, $t$, the input image $x$ and the corresponding text $\mathbf{c}$. We use the denoising U-Net from Stable Diffusion v2 [66] to initialize is the diffusion model. Specifically, given a text caption $\textbf{c}=\{c_{1},c_{2},...\}$, we first encode it through a frozen CLIP text encoder, then feed it into the diffusion model. Given an image $x$, we encode it using the same VAE encoder $E(x)$ that encoded the depth map, concatenated with the depth latent $\mathbf{z}_{t}$, then feed it into the diffusion model.
Training objective. Our training objective is derived using the reparameterization trick and variational bounds, which involves predicting the noise added at each step. The loss function is:
$\mathcal{L}(\theta)=\mathbb{E}_{y,\epsilon,t}\left[\|\epsilon-\epsilon_{\theta}(\mathbf{z}_{t},t,x,\mathbf{c})\|^{2}\right],$
where: $\epsilon\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ is the noise sampled during the forward process, and $\mathbf{z}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{z_{y}}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$ is the noisy depth map at step $t$, where $\bar{\alpha}_{t}=\prod_{s=1}^{t}(1-\beta_{s})$.
[…]
Table 1: Quantitative comparison. PriorDiffusion achieves performance on par with other state-of-the-art affine-invariant depth estimators that use significantly more training images, and outperforms methods trained with a similar number of images. Metrics are reported as percentages. * indicates the results of the model that we re-implemented. Note that our method converges faster compared with Marigold, and is even better with only two-thirds iterations, attributing the language prior which provides a stronger constraint for the fast convergence of the diffusion trajectory. For fair comparison, PriorDiffusion and Marigold are evaluated without ensemble.
Inference. The inference phase begins by sampling a purely noisy latent variable $\mathbf{z}_{T}$ from a standard Gaussian distribution, $\mathbf{z}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. This initial noisy sample represents the starting point of the reverse diffusion process. The goal is to progressively refine this sample through a series of denoising steps until the final depth latent $\mathbf{z}_{0}$ is obtained. Each denoising step is to predict the noise component $\epsilon_{\theta}(\mathbf{z}_{t},t,x,\mathbf{c})$ that needs to be removed from the latent variable at time step $t$. The prediction is made by the diffusion model conditioned on the noisy latent $\mathbf{z}_{t}$, the current timestep $t$, the input image $x$, and the corresponding text description $\mathbf{c}$. The iterative denoising process follows the reverse transition:
$\mathbf{z}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{z}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(\mathbf{z}_{t},t,x,\mathbf{c})\right),$
where $\alpha_{t}$ and $\bar{\alpha}_{t}$ are derived from the noise schedule $\beta_{t}$. This reverse diffusion continues until $t=0$, at which point the fully denoised latent $\mathbf{z}_{0}$ is reached. The predicted depth map $\hat{y}$ is obtained by decoding $\mathbf{z}_{0}$ back into the image space using the pre-trained and frozen VAE decoder:$\hat{y}=D(\mathbf{z}_{0}).$
Figure 6: Controllable depth estimation with language. Each time, we append the newly added sentence to the end of the original sentence and repeat the inference process. Is shows that by providing more details in the language, our method predicts more accurate depth for those specified regions or objects.
4 Experiments
Training datasets. We adopt the training scheme of Marigold [38], and train our PriorDiffusion on two synthetic datasets that cover both indoor and outdoor scenes. The first dataset, HyperSim [65], is a photorealistic collection of 461 indoor scenes. From this, we use the official split to select around 54K samples from 365 scenes, excluding any incomplete samples. RGB images and depth maps are resized to a resolution of 480 × 640 pixels, and the original distance measurements, relative to the focal point, are converted to conventional depth values relative to the focal plane. The second dataset, Virtual KITTI [4, 20], includes synthetic street scenes from 5 settings with varied conditions, such as different weather and camera perspectives. We use the Virtual KITTI 2 [4] version. For training, we select four scenes, totaling approximately 20K samples, and crop the images to align with the KITTI benchmark resolution [25], setting the maximum depth to 80 meters.
Depth normalization for training. Following [63, 90, 38, 62], to standardize the ground truth depth maps $\mathbf{d}$, we apply a linear normalization ensuring that the depth values primarily fall within the range $[-1,1]$, aligning with the input value range of the VAE. It ensures an affine-invariant depth representation that remains robust to data statistics, guaranteeing that all scenes are constrained by near and far planes with extreme depth values. The affine transformation for normalization is defined as:
$\tilde{y*}=\left(\frac{y^{*}-y_{2}}{y_{98}-y_{2}}-0.5\right)\times 2,$ (1)
where $y_{2}$ and $y_{98}$ represent the 2% and 98% percentiles of the depth maps, respectively.
Evaluation datasets. We assess the performance of our PriorDiffusion using 4 real-world datasets that were not part of its training data, following the configuration as Marigold [38]. Both NYUv2 [72] and ScanNet [10] consist of indoor scenes captured with an RGB-D Kinect sensor. For NYUv2, we use the specified test set, containing 654 images. With ScanNet, we randomly select 800 images from the 312 official validation scenes for evaluation. The KITTI dataset [25, 26] features outdoor street scenes with sparse metric depth captured by a LiDAR sensor, where we apply the Eigen test split [14] consisting of 652 images, following the evaluation protocol of [15]. For ETH3D [71], which are also based on LiDAR measurements, we use all 454 samples that include ground truth depth maps. Following the affine-invariant depth evaluation protocol [99, 63, 90, 38, 62], we evaluate first-order threshold accuracy $\delta_{1}$ and mean absolute relative error $AbsRel$. Details of the metric are provided in the supplementary material.
Obtain language description. To implement our method, we require human-provided text descriptions for each image. Since standard benchmarks do not include such descriptions, we propose using off-the-shelf models to generate text descriptions to simulate those a human would provide. For this purpose, we use the visual question-answering model LLaVA v1.6 [49] to generate 10 text descriptions for each training image using different prompts, and 1 text description for each testing image for evaluation. The details of prompt design are shown in the supplementary material. In each training step, we randomly select 1 text out of 10 that aligns with the input image.
Implementation details. We implemented our method using PyTorch, employing Stable Diffusion v2 [66] as the backbone and adhering to the original pre-training setup with the $v$-objective [68]. During training, we used the DDPM noise scheduler [33] with 1000 diffusion steps, while at inference time, the DDIM scheduler [74] was applied with 50 sampling steps. Our training process spanned 30000 iterations with a batch size of 32, achieved through gradient accumulation over 16 steps (batch size of 2 per step) to fit a single GPU. We optimized the model using the Adam optimizer with a learning rate of $3\cdot 10^{-5}$ and included random horizontal flipping as data augmentation.
Quantitative comparison. Shown in Table 3, PriorDiffusion matches the performance of state-of-the-art affine-invariant depth estimators that require significantly more training images, while outperforming methods trained with a similar number of images. Another significant advantage of PriorDiffusion is its faster convergence. Compared to Marigold, even at 20,000 iterations, PriorDiffusion achieves performance metrics comparable to or better than Marigold’s results at 30,000 iterations. This is also demonstrated in Figure 7. This can be attributed to the language prior integrated into our model, which provides additional semantic and geometric prior serving as an extra constrain that speed up diffusion training. An intuitive explanation is that the model expends less effort in learning the geometric properties and spatial relationships of objects from the image, which is high-dimensional and computationally intensive to process. Instead, much of these features are already embedded in the language input, where these features are represented in a condensed and low-dimensional format, which is easier to learn.
Qualitative comparison. Besides the overall quantitative improvement, we claim that language prior can also guide the model’s attention to better perceive the region that is specified in the language description, which aligns more with users’ interests. The visualization for NYUv2 [72] is shown in Figure 4, where our PriorDiffusion achieves more accurate depth prediction for a given input image compared to Margold, especially for instances highlighted in the language description. For instance, in the first row, the soap dispenser is barely distinguishable in the input image due to its color and texture blending with the background. Vision-based depth estimators like Marigold struggle to identify it, resulting in a depth prediction that matches the background. In contrast, the language description explicitly includes the term “a soap dispenser”, providing the model with a geometric and spatial prior that indicates the existence of a soap dispenser and what it should look like. Consequently, our PriorDiffusion can more accurately perceive the depth of the soap dispenser. Similarly, as shown in Figure 5, we visualize results for the KITTI [25] dataset, where objects such as a parked car or a sign are barely discernible from the visual signal alone. As shown in Figure 6, by incorporating more detailed language descriptions, depth predictions are enhanced for the specified regions or objects.
Faster training convergence. One of the key advantages of PriorDiffusion is its faster convergence compared to Marigold. As shown in Figure 7 here, the improved convergence can be attributed to the integration of a language prior in our model, which provides additional semantic and geometric constraints that help accelerate the diffusion process. Due to computation overhead, following the same configuration of Marigold, we use a subset of KITTI and NYUv2 for training-time evaluation.
Figure 7: Convergence speed comparison. PriorDiffusion converges faster during training compared with Marigold, since it’s more efficient to learn 3D property from condensed, low-dimensional language feature.
Figure 8: Performance under different denoising steps. PriorDiffusion consistently outperforms Marigold across various denoising steps, with a faster convergence.
Fewer denoising steps. As demonstrated in Figure 8, and detailed in Table 4 in the supplementary material, we evaluate PriorDiffusion and Marigold with different denoising steps during inference. PriorDiffusion consistently outperforms Marigold across different denoising steps. Remarkably, PriorDiffusion converges in just 10 steps, while Marigold takes 25 steps to achieve convergence. This indicates that the language prior serves as an effective initialization of the diffusion trajectory, enabling faster convergence.
Better perception in small areas. To demonstrate a better perception for small objects and areas mentioned in text, as shown in Figure 4. We use MaskDINO to obtain panoptic segmentation masks for all images in NYUv2 test set, then evaluate the performance for masks that occupied less than 5%, 10%, and 20% of an image. Table 2 shows that small areas are harder to perceive compared with large areas, and PriorDiffusion obtains better improvement over small areas, as language prior provides geometric and semantic properties that guide the diffusion model to better perceive small areas and tiny structures.
NYUv2 Standard Small Area (5%) Small Area (10%) Small Area (20%) $\delta_{1}$ AbsRel↓ $\delta_{1}$ AbsRel↓ $\delta_{1}$ AbsRel↓ $\delta_{1}$ AbsRel↓ Marigold 95.7 6.1 91.7 9.0 92.2 8.4 93.9 7.3 Ours 95.9 5.9 92.8 8.3 93.1 7.9 94.6 6.8
Table 2: Better perception in small areas. We evaluate the performance for panoptic segmentation masks that occupied less than 5%, 10%, and 20% of an image. PriorDiffusion obtains better improvement over small areas due to geometric properties of those areas offered by the language prior.
Figure 9: Failure cases. Incorrect text (here we replace ”a bookshelf with glass” with ”a window with curtains”) misguides PriorDiffusion to perceive the geometry of objects.
5 Discussion
Conclusion. We present PriorDiffusion, a novel approach that integrates language prior with diffusion models to enhance monocular depth estimation. By leveraging geometric prior learned by text-to-image model that associates with language, PriorDiffusion overcomes traditional ambiguities and visual nuisances associated with monocular depth estimation, achieving state-of-the-art performance across various zero-shot benchmarks, while maintaining a faster convergence for training and inference diffusion trajectory.
Limitation. One primary limitation is its dependence on the accuracy and detail of the provided language descriptions. Shown in Figure 9, ambiguous or misleading text inputs from human users can result in suboptimal depth predictions and potentially compromise the model’s performance. Last, our method requires language input from users, which may not always be feasible. In cases where user-provided descriptions are unavailable, as demonstrated in Figure 5 in the supplementary material, our method can still function with standardized template prompts. Another potential solution is to generate language descriptions on the fly using a vision-language model [11], an image captioner [34], or by following methods like RSA [99] to create structured text from the output of visual algorithm [116, 42]. While we have not explored this approach in this paper, it presents a promising avenue for future research.
Supplementary Material
A. Dataset Details
We train our model on two synthetic datasets, Hypersim [65] and Virtual Kitti [4], and conduct zero-shot evaluations on four additional real-world datasets that were not part of its training data, NYUv2 [72], KITTI [25], ScanNet [10], and ETH3D [71]. Details of each dataset are provided below.
A.1. Training Datasets
Hypersim [65] is a photorealistic synthetic dataset designed for comprehensive indoor scene understanding, and is introduced since obtaining per-pixel ground truth labels from real images is often challenging or impossible for many essential scene understanding tasks. This dataset is created using a vast collection of synthetic scenes developed by professional artists, resulting in 77,400 images across 461 indoor scenes with detailed per-pixel annotations and corresponding ground truth geometry. HyperSim is built exclusively using publicly accessible 3D assets. It includes complete scene geometry, material properties, and lighting information for each scene. Also, it provides dense per-pixel semantic instance segmentations and comprehensive camera details for each image. Further, it decomposes each image into diffuse reflectance, diffuse illumination, and a non-diffuse residual component that captures view-dependent lighting effects. In terms of training split, as mentioned in the Experiments Section, we utilize the official dataset split to select approximately 54,000 samples from 365 scenes, ensuring that any incomplete samples are excluded. The RGB images and depth maps are resized to a resolution of 480 × 640 pixels, and the original distance measurements, defined relative to the focal point, are transformed into standard depth values relative to the focal plane.
Virtual Kitti [4, 20] Virtual KITTI is a photorealistic synthetic video dataset created for training and evaluating computer vision models on various video understanding tasks, including object detection, multi-object tracking, scene-level and instance-level semantic segmentation, optical flow, and depth estimation. The dataset comprises 50 high-resolution monocular videos (a total of 21,260 frames) generated from five distinct virtual urban environments, each presented under varying imaging and weather conditions. These virtual scenes were developed using the Unity game engine and an innovative real-to-virtual cloning technique. The synthetic videos come with precise, fully automatic annotations for 2D and 3D multi-object tracking, as well as per-pixel category, instance, flow, and depth labels. We use its upgraded version, Virtual KITTI 2 [4], which consists of the same five sequence clones as Virtual KITTI, with increased photorealism. It takes advantage of recent advancements in lighting and post-processing within the game engine, making the variations in the virtual sequences more closely mimic real-world changes in conditions. For training, we choose four scenes, comprising around 20,000 samples, and crop the images to match the resolution of the KITTI benchmark [25]. The maximum depth is capped at 80 meters.
A.2. Evaluation Datasets
NYUv2 [72] dataset comprises 24,231 synchronized RGB images and depth maps at a resolution of $640\times 480$, representing various indoor scenes such as homes, offices, and commercial spaces, captured using a Microsoft Kinect. The standard split includes 249 training scenes and 215 test scenes. For our experiments, we use the official test set. Consistent with prior works [2, 97, 98, 99, 38], we exclude samples without valid ground truth, resulting in 654 valid images for evaluation. We perform evaluation on NYUv2 over a depth range spanning from $1\times 10^{-3}$ to 10 meters.
KITTI [25, 26] contains 61 driving scenes with research in autonomous driving and computer vision. It contains calibrated RGB images with synchronized point clouds from Velodyne lidar, inertial, GPS information, etc. Following prior works [2, 97, 98, 99, 38], we used Eigen split [14]. It consists of 652 testing images after filtering out images without valid ground truth. We follow the evaluation protocol of [15] for our experiments.
ScanNet [10] is an extensive RGB-D video dataset containing 2.5 million views in more than 1500 scans, annotated with 3D camera poses, surface reconstructions, and instance-level semantic segmentations. Data was collected using an RGB-D capture system (with a Kinect sensor) that includes automated surface reconstruction and crowdsourced semantic annotation. We use the same evaluation configuration of Marigold [38], where 800 images are randomly selected from the 312 official validation scenes for evaluation.
ETH3D [71] is a multi-view stereo and 3D reconstruction benchmark encompassing a diverse range of indoor and outdoor scenes. High-precision laser scanning was used to obtain the ground truth geometry. Images were captured using both a DSLR camera and a synchronized multi-camera rig with varying fields-of-view. For evaluation, following Marigold [38], we use all 454 samples that include ground truth depth maps.
B. Implementation Details
B.1. Visualization
When visualizing the ground truth depth map, we apply the same affine transformation we used in training. As described in the Experiments Section, we apply a linear normalization ensuring that the depth values primarily fall within the range $[-1,1]$. The affine transformation for normalization is defined as:
$\tilde{y*}=\left(\frac{y^{*}-y_{2}}{y_{2}-y_{98}}-0.5\right)\times 2,$ (2)
where $y_{2}$ and $y_{98}$ represent the 2% and 98% percentiles of the depth maps, respectively. Then, we apply min-max normalization to both the ground truth and predicted depth maps, scaling them to integer values within the range [0, 255]. These normalized depth maps are then visualized using the OpenCV MAGMA colormap.
It is important to note that, unlike Marigold [38], which applies linear fitting to the ground truth for error correction, we do not use this approach. While linear fitting can adjust predictions to more closely align with the ground truth, it does not accurately reflect the true distribution of the depth map predictions or provide a clear assessment of prediction quality. Instead, we enhance visualization by applying the same training normalization to the zero-shot evaluation dataset, avoiding linear fitting and its error correction effects, resulting in more authentic and insightful visualization outcomes.
B.2. Training details.
We implemented our method using PyTorch, employing Stable Diffusion v2 [66] as the backbone and maintaining the original pre-training configuration with the $v$-objective [68]. The training process utilized the DDPM noise scheduler [33] with 1,000 diffusion steps, while at inference time, the DDIM scheduler [74] was employed with 50 sampling steps for faster results. Our training setup spanned 30,000 iterations, with an effective batch size of 32 achieved through gradient accumulation over 16 steps (with a per-step batch size of 2) to fit on a single Nvidia RTX 3090 GPU. We used the Adam optimizer with a learning rate set at $3\cdot 10^{-5}$ and included random horizontal flipping with a probability of 0.5 as data augmentation. For depth normalization, we employed a scale and shift-invariant method with clipping enabled and set the normalization range between -1.0 and 1.0, using a 0.02 min-max quantile to maintain robustness. This normalization strategy was applied during training and zero-shot evaluations for consistency. The training noise scheduler initialized from the pre-trained Stable Diffusion v2 [66] model maintained a noise strength of 0.9 and incorporated an annealed strategy to progressively reduce noise levels. We saved checkpoints every 50 iterations, with backup, validation, and visualization checkpoints set at intervals of 2,000 iterations. The training process typically converged after approximately 20,000 iterations, though we extended training to 30,000 iterations for thorough coverage. We used mean squared error (MSE) as the loss function, with reduction set to “mean” for averaged loss calculation. A customized iteration-wise exponential scheduler is applied which adjusts the learning rate iteratively using an exponential decay function. It decays the learning rate to 1% of its initial value over 25,000 iterations with a warmup phase of 100 steps. For text generation, generating a single caption for an image using LLaVA v1.6 on an RTX 3090 takes approximately 3.6 seconds, and we generate 10 captions for each image. For training, we generate 740,000 captions, using 740 GPU hours on an RTX 3090.
B.3. Evaluation metric
Following the affine-invariant depth evaluation protocol [99, 63, 90, 38, 62], for each image and the predicted relative depth $y$, we fit a pair of scalars denoting the scale and shift parameters of the transformation: $(\hat{\alpha},\hat{\beta})=g_{\psi}(y,y^{*})\in\mathbb{R}^{2}$. The metric depth prediction is obtained by $\hat{y}=\hat{\alpha}\cdot y+\hat{\beta}$ such that:
$\psi^{*}=\arg\min_{\psi}\frac{1}{|{M}|}\sum_{(i,j)\in\Omega}M(i,j)|\hat{y}(i,j)-y(i,j)|$ (3)
where $\hat{y}=\hat{\alpha}\cdot y+\hat{\beta}$ denotes the predicted metric-scale depth aligned from relative depth $y$, $(i,j)\in\Omega$ denotes an image coordinate, and $M:\Omega\mapsto\{0,1\}$ denotes a binary mask indicating valid coordinates in the ground truth depth $y^{*}$ with values greater than zero. Then, we follow [6, 48, 106, 98, 99] to evaluate using first-order threshold accuracy, calculated as:
$\delta_{1}=\%\text{ of }y(i,j)\text{ s.t. }\max(\frac{{y}(i,j)}{y^{*}(i,j)},\frac{y^{*}(i,j)}{{y}(i,j)})<1.25$ (4)
and mean absolute relative error, calculated as:
$AbsRel=\frac{1}{|{M}|}\sum_{(i,j)\in\Omega}\frac{|y^{*}(i,j)-{y}(i,j)|}{y^{*}(i,j)}$ (5)
C. Additional Experiments
C.1. Additional visualizations
We provide additional visualization and analysis for indoor scenes in Figure 10 and outdoor scenes in Figure 11. We use several samples in the NYUv2 [72] and KITTI [25, 26] dataset across diverse types of scenes. We have provided examples in captions under each figure. These visualizations demonstrate that leveraging the language prior within the text-to-image diffusion model enhances the model’s ability to understand the geometric characteristics of the specified regions and objects. It shows that language plays a critical role in guiding the model’s attention to relevant regions and providing context for improved depth prediction. It highlights subtle or easily overlooked details, such as small objects or instances, and enhances the perception of complex scenes with multiple objects or intricate surfaces. Additionally, language descriptions offer an essential context for partially observed or occluded objects, enabling the model to infer details that visual cues alone might miss. By integrating language prior, the model achieves a more comprehensive and accurate understanding of scenes, especially in challenging scenarios.
Figure 10: Additional visualization on NYUv2. Compared to Marigold, our PriorDiffusion demonstrates better depth prediction, particularly for instances specified in the language description (highlighted in red text and marked with red boxes). The language description effectively guides the model’s attention to relevant regions, especially those easily overlooked by visual cues due to a small size or a transparent texture, such as “a green spray can” in the 1st row and “two crystal ornaments” in the last row. It also improves perception under challenging visual conditions, like “shelves” in the 4th row and “a white dresser” in the 5th row, both of which are under poor illumination and are difficult to tell from visual solely. Additionally, it supports complex reasoning about scene layouts that might be misinterpreted from visual cues alone, such as “a doorway in the background” in the 6th row. Furthermore, it provides critical context for partially observed or occluded objects, such as “a partially open door revealing a storage area” in the 2nd row and “a small bar cart in the corner” in the 7th row.
Figure 11: Additional visualization on KITTI. Compared to Marigold, our PriorDiffusion demonstrates superior depth prediction, particularly for instances specified in the language description (highlighted in red text and marked with red boxes). The language description effectively guides the model’s attention to relevant regions that might otherwise be overlooked due to their small size or subtle visual cues. Examples include “banners advertising a store” in the upper 1st column and “a red car and a blue car” in the upper 2nd column. Additionally, it enhances perception in complex scenes featuring multiple objects or intricate surfaces. For instance, it accurately captures “surrounded by trees” in the lower 1st column, “dense greenery and tall trees” in the upper 3rd column, and “bicycles, and people walking and interacting” in the lower 3rd column. Furthermore, the language descriptions provide essential context for partially observed or occluded objects, such as “buildings covered by trees” in the lower 2nd column.
Method NYUv2 KITTI ETH3D $\delta_{1}$ AbsRel↓ $\delta_{1}$ AbsRel↓ $\delta_{1}$ AbsRel↓ “An image” 95.7 6.1 89.8 10.7 95.1 6.8 Template A 95.8 6.0 90.3 10.6 95.5 6.4 Template B 95.7 6.1 90.5 10.7 95.6 6.5 Template C 95.9 6.0 90.2 10.6 95.3 6.6 Template D 95.6 5.8 90.4 10.5 95.4 6.4 Template E 95.8 5.9 90.3 10.7 95.6 6.7 Template F 95.9 6.0 90.5 10.5 95.4 6.6 Template G 95.8 6.1 90.3 10.6 95.5 6.3 Template H 95.7 5.9 90.2 10.7 95.7 6.3 Template I 95.8 6.1 90.6 10.5 95.4 6.6 Template J 95.8 6.0 90.5 10.6 95.6 6.4 Template K 95.7 5.8 90.4 10.5 95.5 6.3 Ours 95.9 5.9 90.6 10.4 95.7 6.5
  • Template A: “Describe the image in one sentence. Explain the image by identifying key objects and their distances from the viewpoint, noting any perspective lines or depth cues that indicate the three-dimensional structure of the scene.”
  • Template B: “Describe the image in one sentence. Describe the image by detailing the foreground, midground, and background objects, emphasizing their relative distances and spatial positioning within the scene.”
  • Template C: “Describe the image in one sentence. Describe the image by detailing the foreground, midground, and background objects, emphasizing their relative distances and spatial positioning within the scene.”
  • Template D: “Describe the image in one sentence. Provide an in-depth description of the image, focusing on the scale and depth of each visible object and how they overlap or are spaced from one another.”
  • Template E: “Describe the image in one sentence. Analyze the image by discussing the size and arrangement of objects, their positions relative to one another, and any changes in texture or clarity that indicate varying depths across the scene.”
  • Template F: “Describe the image in one sentence. Describe the scene with attention to depth, specifying which elements appear closer or farther from the viewer and how shadows or lighting contribute to the perception of depth.”
  • Template G: “Describe the image in one sentence. Highlight the depth relationships in the image by describing which objects are in the foreground, which are in the background, and how their relative sizes help convey distance.”
  • Template H: “Describe the image in one sentence. Focus on any natural or man-made structures in the image and describe how their orientation and placement give a sense of depth or perspective.”
  • Template I: “Describe the image in one sentence. Describe how elements like roads, pathways, or fences create leading lines that guide the viewer’s eye into the depth of the scene.”
  • Template J: “Describe the image in one sentence. Explain how differences in lighting or shadowing in the image indicate which parts are nearer or further away from the observer.”
  • Template K: “Describe the image in one sentence. Analyze the spatial arrangement of the main objects and describe any overlapping or occlusion that suggests depth relationships between them. ”
  • Ours: “Describe the image in one sentence, assuming it’s a real-world image, pay more attention to objects, their spatial relationships, and the overall layout.”
Table 3: Ablation for prompts to generate language description. Prompts are used to prompt LLaVA to generate language descriptions for each image. While the performances among different prompts may vary, they remain consistently comparable, as long as they are meaningful and mimic human descriptions.
C.2. Ablation for prompts to generate text
To study the effect of different prompts for text generation, here we use the visual question-answering model LLaVA v1.6 Mistral [49] to generate one text for each training image and testing image. The prompt we use should elicit responses that capture essential details, including the positioning of objects, their interactions, and notable features that influence monocular depth estimation. We generate different prompts using ChatGPT 4o [54], with the prompt:
“Generate prompt for a vision-language model to generate language description for each given image in one sentence. The prompt we use needs to elicit responses that include essential details, such as the positioning of objects, their interactions, and notable features that may impact depth estimation.”
We present the results in Table 3, showcasing various language descriptions generated by LLaVA under different prompts. While the performances exhibit some variation, they remain consistently comparable. This demonstrates that as long as meaningful language descriptions of 3D scenes, resembling natural descriptions provided by humans, are provided, PriorDiffusion can effectively leverage the language prior to enhance monocular depth estimation.
C.3. Different denoising steps.
As demonstrated in Figure 8, we evaluate PriorDiffusion and Marigold with different denoising steps during inference. The extracted performances are shown in Table 4. With more denoising steps, performance gradually improves. PriorDiffusion consistently outperforms Marigold across various denoising steps, converging in just 10 steps, while Marigold requires 25 steps. This suggests that the language prior could offer a good initialization, speeding up the denoising process and accelerating convergence.
Steps NYUv2 KITTI ETH3D ScanNet $\delta_{1}$ AbsRel↓ $\delta_{1}$ AbsRel↓ $\delta_{1}$ AbsRel↓ $\delta_{1}$ AbsRel↓ Marigold 1 48.8 33.8 25.2 59.1 50.1 37.5 60.3 25.7 2 78.7 16.9 72.2 18.0 86.0 12.7 72.1 19.2 3 92.6 8.2 77.7 15.3 92.0 8.6 87.1 11.4 4 94.2 7.0 80.4 14.1 93.6 7.6 90.7 9.3 5 94.7 6.7 82.4 13.4 94.3 7.3 91.9 8.6 10 95.1 6.4 86.5 11.9 94.9 6.9 93.2 7.8 15 95.1 6.4 87.4 11.6 95.0 6.9 93.4 7.7 20 95.2 6.3 88.0 11.4 95.0 6.9 93.7 7.5 25 95.3 6.3 88.2 11.3 95.0 6.9 93.7 7.5 50 95.3 6.3 88.5 11.2 95.0 6.9 93.8 7.5 PriorDiffusion 1 48.8 33.8 25.2 59.1 50.1 37.5 60.3 25.7 2 83.2 14.1 74.7 16.9 86.5 12.0 75.0 17.8 3 94.3 7.1 81.2 13.9 93.0 7.7 90.6 9.5 4 95.5 6.3 84.7 12.4 94.7 6.9 93.5 7.8 5 95.7 6.0 87.0 11.6 95.3 6.6 91.9 8.6 10 96.0 5.9 89.9 10.5 95.7 6.5 94.9 6.8 15 96.0 5.9 90.1 10.4 95.8 6.5 94.9 6.8 20 96.0 5.9 90.3 10.4 95.8 6.5 94.9 6.7 25 96.0 5.9 90.3 10.4 95.8 6.5 94.9 6.7 50 96.0 5.9 90.3 10.4 95.8 6.5 94.9 6.7
Table 4: Performance for different denoising steps. PriorDiffusion consistently outperforms Marigold across various denoising steps, with a significantly faster convergence speed for the diffusion process.
C.4. Template prompt comparison.
In situations where user-provided descriptions are not available, we aim to explore whether our method can still perform effectively using either a blank input or standardized template prompts. If no specific language description that aligns with the scene is given by the user, we could instead use some pre-defined template prompt, such as “An image” or more elaborate prompts like, “A complex 3D scene with varying objects at different distances.”, as input of PriorDiffusion to maintain its performance. ” These templates, as presented in Table 5, assist in preserving the model’s performance when language input is not practical. Consequently, using template prompts as input during inference may differ from the training domain, leading to a significant covariate shift in language descriptions. This issue can be mitigated by incorporating template prompts as part of the input during training, enabling PriorDiffusion to better adapt to various scenarios where meaningful language descriptions are not feasible. Nevertheless, PriorDiffusion can achieve comparable or even better performance with Marigold when language input is not feasible, thus using a fixed prompt instead, showing that simply training with language can improve the performance of the depth estimator.
Method NYUv2 KITTI ETH3D $\delta_{1}$ AbsRel↓ $\delta_{1}$ AbsRel↓ $\delta_{1}$ AbsRel↓ Blank text input 95.5 6.2 89.3 10.9 95.0 6.9 “An image” 95.7 6.1 89.8 10.7 95.1 6.8 Template A 95.8 6.0 89.9 10.7 95.3 6.8 Template B 95.7 6.1 89.8 10.7 95.2 6.8 Template C 95.8 6.1 89.8 10.7 95.3 6.8 Marigold 95.7 6.1 89.7 10.7 95.4 6.9 With text input 96.0 5.9 90.3 10.4 95.8 6.5
  • a
    Template A: ‘A complex 3D scene with varing objects at different distances.”
  • b
    Template B: ‘A structured environment with intricate patterns and designs that create depth and guide the eye through various focal points.”
  • c
    Template C: ‘An elaborate scene with overlapping objects that create a sense of distance and spatial hierarchy within the environment.”
Table 5: Template prompt comparison. Metrics are reported as percentages. PriorDiffusion can achieve comparable or even better performance with Marigold when language input is not feasible, thus using a fixed prompt instead. The last row is the original PriorDiffusion with text input.
C.5. Obtain language description.
To generate text descriptions for images, for training images, we use two different version of LLaVA v1.6 [49], Mistral and Vicuna, each with 5 different prompts, to generate 10 text descriptions for each training image:
  • “Describe the image in one sentence, assuming it’s a real-world image.”
  • “Provide a one-sentence description of the image, pay attention to object type, assuming it’s a real-world image.”
  • “Capture the essence of the image in a single sentence, pay attention to object relationship, assuming it’s a real-world image.”
  • “Condense the image description into one sentence, pay attention to object size, assuming it’s a real-world image.”
  • “Express the image in just one sentence, pay attention to the overall layout, assuming it’s a real-world image.”
For testing images, we use LLaVA v1.6 Mistral, then prompt this model with:
  • “Describe the image in one sentence, assuming it’s a real-world image. Pay close attention to objects, their spatial relationships, and the overall layout.”
This prompt is designed to elicit responses that include essential details, such as the positioning of objects, their interactions, and notable features that may impact depth estimation. Note that all the training data we use is synthetic data, and by emphasizing “assuming it’s a real-world image,” we ensure that the descriptions align with the types of inputs and scenarios the model will encounter.